Operations Report – March 16th

Crew 206 Operations Report 16 – MAR – 2019

SOL: 20

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 11 hours

Solar— SOC % – Turned off at 77%, Still off at 75% (1900)

Diesel Reading – 55%
Propane Reading – 69%

Ethanol Free Gasoline – 2 gallons

Water (Auxiliary tank) – Not in use

Water (Static tank) – 280 gallons

Water differential (static tank) – 30 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 193 gallons

Static to Loft Pump used – Yes

Water Meter: 0141813.6 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 54.5 hours

Beginning charge: 100%

Ending charge: 46% (48% currently)

Currently charging: Yes

Curiosity rover used: Used

Hours: 82.7 hours

Beginning charge: 100%

Ending charge: 63%

Currently charging: Yes

Spirit rover used: Used

Hours: 83.3 hours

Beginning charge: 100%

Ending charge: 40% (47% currently)

Currently charging: Yes

Notes on rovers: Spirit and Opportunity weren’t charging due to a socket failure. It has been fixed.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff and Crew 206 for water refilling.

Staff Crew Car used and why, where? Used by Crew 207 to come to MDRS.

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Nothing more to report.